Use the sign of >,
 (-3) + 7 – (19) [ ] 15 – 8 + (-9)

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Let us take Left Hand Side (LHS) = (-3) + 7 - 19
= -3 + 7 - 19
= -22 + 7
= -15
Now, Right Hand Side (RHS) = 15–8 + (-9)
= 15–8 - 9
= 15–17
= -2
By comparing LHS and RHS
LHS < RHS
-15 < -2
∴(-3) + 7–(19) [<] 15–8 + (-9)

Thanks
Manika gupta
3 years ago
Dear Student
You can see your solution here.

LHS
(-3) + 7 – (19) = -15
RHS
15 – 8 + (-9)=-2

And we know that -2 is greater than -15.

AskIITians expert.